Status Report

NASA ISS On-Orbit Status 5 February 2009

By SpaceRef Editor
February 5, 2009
Filed under , , ,
NASA ISS On-Orbit Status 5 February 2009
http://images.spaceref.com/news/iss.69.jpg

All ISS systems continue to function nominally, except those noted previously or below.

In further preparation of Progress 31P undocking tonight (~11:11pm EST), FE-1 Lonchakov & CDR Fincke –

  • Finished packing and tying down trash and excess cargo while logging moves in the IMS (Inventory Management System),
  • Reported on the finished stowage job to TsUP-Moscow,
  • Activated 31P and tore down the ventilation air duct to the DC-1,
  • Worked jointly to remove the SSVP BZV rigidizing quick-disconnect screw clamps,
  • Downlinked the situational/documentary video of the cargo configuration for ground inspection,
  • Closed DC1-to-Progress transfer hatches (~6:10am EST), and
  • Started the usual one-hour hatch leak check on the connecting vestibule.

The CDR performed the scheduled lens change on the EarthKAM system at the Node-2 hatch window, going from 50mm to 180mm focal length. The lens change had to be performed on the day side while EarthKAM was not taking pictures, with Fincke able to see the ground when focusing the camera. [EarthKAM was activated on 2/2 for a new session. This is the 30th time for EarthKAM aboard the ISS and the second time on Increment 18. The payload runs without crew intervention. EK is using a Kodak ESC 460C electronic still camera with 50mm and 180mm lenses, powered by 16Vdc from a 28 Vdc adapter, taking pictures by remote operation from the ground, without crew interaction. It is available for students who submit image requests and conduct geographic research (student participants include over 84 teachers who are being trained to use EK). The requests are uplinked in a camera control file to the ThinkPad A31p laptop which then activates the camera at specified times and receives the digital images from the camera’s storage card on its hard drive, for subsequent downlink via OpsLAN. ]

Fincke & Magnus prepared for the new US payload SPICE (Smoke Point In Co-flow Experiment) by reviewing instructions material and tagging up with ground specialists at POIC (Payload Operation & Integration Center). Sandy then activated the MSG (Microgravity Science Glovebox) via its A31p laptop and readied the equipment for operation. [Setup steps included configuring the MSG facility, unstowing SPICE hardware and installing it into the MSG work volume, then installing & checking out the videocam & digital still camera for documentary imagery, and performing an igniter test. Several hours later, Magnus configured the MSG for Standby and powered down the A31p. The SPICE determines the point at which gas-jet flames (similar to a butane-lighter flame) begin to emit soot (dark carbonaceous particulate formed inside the flame) in microgravity. Studying a soot emitting flame is important in understanding the ability of fires to spread and in control of soot in practical combustion systems space.]

Sandy also turned off and disassembled the MPC (Multi-Purpose Converter) HD video system which she had activated yesterday for some troubleshooting by the ground. [The HD MPC signal has been exhibiting significant glitches and dropouts during recent PAO events. Until the issue is solved, the crew was requested to set up standard definition (SD) television simultaneous with HD MPC, and to conduct the events with audio on S/G 2 since the audio also can drop out in certain circumstances.]

The CDR performed the regular periodic US WRS (Water Recovery System) sampling. [After first setting up the TOCA (Total Organic Carbon Analyzer) by connecting its power cable to the UOP1 J3 outlet panel and hooking up its primed WPA (Water Processing Assembly) water sample hose to the WRS Rack 1, Sandy collected samples from the PWD (Potable Water Dispenser) Hot needle outlet for subsequent inflight processing with the TOCA plus WMK (Water Microbiology Kit) with MCD (Microbial Capture Device) and CDB (Coliform Detection Bag). After the analyses, the usual water reclamation from the sample bags via an absorbing towel (to be dried by airing) and data recording (from TOCA USB drive into the SSC7 laptop) concluded the activities. WRS sampling & checkouts are being conducted for 90 days, i.e., every 4 days: WRS water hose (TOCA inflight analysis) & microbial bag sample (inflight bacterial visual enumeration plus archival for return on 15A), every 8 days: an archival water sample (return on 15A), and monthly: a TOCA bag sample from PWD (tested inflight).]

In preparation for a new run with the Russian payload TKhN-7 SVS (Self-Propagating High-Temperature Synthesis), Yuri performed an end-to-end checkout of the SVS payload equipment and camera gear. [SVS uses its own camera, “Telescience” hardware from PK-3 (Plasma Crystallization) and the onboard Klest TV system for researching self-propagating high-temperature fusion of samples in space.]

Sandy Magnus conducted the weekly 10-min. CWC (Contingency Water Container) inventory as part of on-going WRM (Water Recovery & Management) assessment of onboard water supplies. Updated “cue cards” based on the crew’s water calldowns are sent up every other week. [The new card (18-0006K) lists 41 CWCs (~1,245.4 L total) for the four types of water identified on board: technical water (665.3 L, for Elektron electrolysis, incl. 110.6 L currently off-limits, filled from WPA and pending sample analysis on the ground), potable water (530.4 L, incl. 174.6 L currently off-limit because of Wautersia bacteria), condensate water (0.0 L), waste/EMU dump and other (49.7 L, including the new CWC-I with 3 L). Wautersia bacteria are typical water-borne microorganisms that have been seen previously in ISS water sources. These isolates pose no threat to human health.]

In preparation of the 15A spacewalks, the CDR & FE-2 spent ~30 min in the US A/L (Airlock) to check out two SAFER (Simplified Aid for EVA Rescue) units.

On the FCF (Fluids & Combustion Facility) in the Lab, Sandy removed alignment guides from the CIR (Combustion Integrated Rack) to allow its PaRIS (Passive Rack Isolation System) to be activated before FCF operations begin.

The FE-1 performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers, replacement of EDV-SV waste water and EDV-U urine containers and performing US condensate processing (transfer from CWC to EDV containers) if condensate is available.]

Yuri also conducted the regular daily IMS (Inventory Management System) maintenance task by updating/editing the IMS standard “delta file” including stowage locations for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Mike Fincke had ~45 min to video another set of onboard scenes, today focusing on the state of the “Quest” A/L (EVA tools, EMU configs, generic layout of METOX, etc.) for the benefit of the STS-119/15A crew.

The station residents completed their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR, FE-1, FE-2), ARED advanced resistive exerciser (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

At ~8:25am, the three crewmembers conducted a 30-min teleconference with members of the STS-119 to “compare notes” and discuss transfer details during the docked period.

At 11:50am, Mike & Sandy set up the G1 video camcorder with MPC (Multipurpose Converter) and IPU (Image Processing Unit) as well as standard definition (SD) television and at ~12:00pm supported a 20-min PAO TV event of two interactive interviews of 10 min each, one with Space.com (Tariq Malik), the other WBBM Radio in Chicago, IL (Steve Grzanich).

For his discretionary choice, CDR Fincke had two jobs on his “job jar” task list for today: (a) Installing improvised flywheel handle covers on the ARED, and (b) adjusting (reducing) the thickness of Delrin wear strips on the ARED lift bar slide tracks.

MT Relocation Update: The Mobile Transporter was sucxcessfully moved from WS4 (Worksite 4) to WS1, then to WS6 and checked out for Flight 15A.

Progress 31P Undocking: In preparation for the US-to-RS attitude control handover maneuver tonight at 9:00pm EST for the Progress M-01M/31P undocking at ~11:11pm, Mike closed the protective shutters of the JPM & Lab science windows. They can be opened again tomorrow at ~3:18am EST earliest.

ISS Orbit (as of this morning, 7:36am EST [= epoch]):
Mean altitude — 356.6 km
Apogee height — 362.2 km
Perigee height — 351.1 km
Period — 91.67 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0008209
Solar Beta Angle — -52.9 deg (magnitude increasing)
Orbits per 24-hr. day — 15.71
Mean altitude loss in the last 24 hours — 48 m
Revolutions since FGB/Zarya launch (Nov. 98) — 58524

Significant Events Ahead (all dates Eastern Time, some changes possible!):
02/05/09 — Progress M-01M/31P undocking & deorbit (~11:11pm EST)
02/10/09 — Progress 32P launch
02/19/09? — STS-119/Discovery/15A launch – S6 truss segment (4:41am EST)—“NOT EARLIER THAN”
02/13/09 — Progress 32P docking (2:20am EST); [crew wake: 10:30pm on 2/12]
02/21/09? — STS-119/Discovery/15A docking
03/02/09? — STS-119/Discovery/15A undocking
03/05/09? — STS-119/Discovery/15A landing
03/26/09 — Soyuz TMA-14/18S launch
03/28/09 — Soyuz TMA-14/18S docking (DC1)
04/07/09 — Soyuz TMA-13/17S undocking
04/07/09 — Progress 32P undocking & deorbit
05/12/09 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/27/09 — Soyuz TMA-15/19S launch
Six-person crew on ISS
08/06/09 — STS-128/Discovery/17A – MPLM (P), LMC, last crew rotation
08/XX/09 — Soyuz 5R/MRM2 (Russian Mini Research Module, MIM2) on Soyuz
09/XX/09 — H-IIB (JAXA HTV-1)
11/12/09 — STS-129/Atlantis/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A – MPLM(P), LMC
04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4
12/XX/11– Proton 3R/MLM w/ERA.

SpaceRef staff editor.